For a thermodynamic system, isobaric coefficient of volume expansion (a) and isothermal compressibility (b) are defined as

a=(1/v) (dv/dT)p    b= -(1/v) (dv/dp)T

Show that for an isochoric change, b dp = a dT.


1
Expert's answer
2021-02-04T02:32:58-0500
"\frac{a}{b}=-\frac{(\frac{\partial V}{\partial T})_p}{(\frac{\partial V}{\partial p})_T}\\\frac{a}{b}=-\left(\frac{\partial V}{\partial T}\right)_p\left(\frac{\partial p}{\partial V}\right)_T\\\frac{a}{b}=\left(\frac{\partial p}{\partial T}\right)_V"